PLEASE HELLPPPP AND MAKE SURE IT IS CORECTTTT
In a recent year, the Better Business Bureau settled 75% of complaints they received. (Source: USA Today, March 2. 2009) You have been hired by the Bureau to investigate complaints this year involving computer stores. You plan to select a random sample of complaints to estimate the proportion of complaints the Bureau is able to settle. Assume the population proportion of complaints settled for the computer stores is the 0.75, as mentioned above. Suppose your sample size is 214. What is the probability that the sample proportion will be within 5 percent of the population proportion? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations. Answer- Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Population proportion, p = 0.75

Sample proportion, hat p_1 = 0.75 - 0.05 = 0.70

Sample proportion, hat p_2 = 0.75 + 0.05 = 0.80

Probability that the sample proportion will be within 5 percent of the population proportion =

P(0.70 << 0.80)

0.80 -p 0.70-p

0.70 0.75 0.80- 0.75 0.75(1-0.75120.750-0.757214 V0.75(1- 0.75)/214

P (-1.6892<z< 1.6892)

P(z < 1.6892)-P(z <-1.6892)

0.9544-0.0456-0.9088

Answer = 0.9088